OpenStudy (anonymous):

Solve : 9^x=45 Help ASAP

OpenStudy (mrcoolguy):

x=log45/2log3

OpenStudy (anonymous):

is that the answer or i solve for that

OpenStudy (mrcoolguy):

its the answer

OpenStudy (anonymous):

can you show me the steps?

OpenStudy (anonymous):

@MrCoolGuy

OpenStudy (solomonzelman):

\(\Large \color{#000000 }{ \displaystyle a^{\color{red}{b}} =c}\) \(\Large \color{#000000 }{ \displaystyle \log_a( a^{\color{red}{b}} )=\log_a(c)}\) \(\Large \color{#000000 }{ \displaystyle \color{red}{b}\log_a( a)=\log_a(c)}\) \(\Large \color{#000000 }{ \displaystyle \color{red}{b}=\log_a(c)}\)

OpenStudy (lvana):

I hope this makes it a little easier on the eyes, but according to @SolomonZelman reference, you should be able to apply the rule flawlessly.~

OpenStudy (anonymous):

oh so it is right thank you

OpenStudy (solomonzelman):

If you apply the rules of logarithms\( ; \) ■ Rule 1: \(\color{#000000 }{ \displaystyle \log_x (x)=1}\) (for positive x, except x=1) ■ Rule 2: \(\color{#000000 }{ \displaystyle \log_x (x^y)=y\times \log_x(x) }\) (for all exponents "y", the exponent goes on outside as shown)

OpenStudy (solomonzelman):

If you want examples, proves, or other references of anything, let me know.

OpenStudy (ryandane):

O.O